# Evaluating and Mitigating LLM-as-a-judge Bias in Communication Systems

### v1

@@ Abstract
Large Language Models (LLMs) are increasingly being used to autonomously evaluate the quality of content in communication systems, e.g., to assess responses in telecom customer support chatbots. However, the impartiality of these AI “judges” is not guaranteed, and any biases in their evaluation criteria could skew outcomes and undermine user trust. In this paper, we systematically investigate judgment biases in two LLM-as-a-judge models (i.e., GPT-Judge and JudgeLM) under the point-wise scoring setting, encompassing 11 types of biases that cover both implicit and explicit forms. We observed that state-of-the-art LLM judges demonstrate robustness to biased inputs, generally assigning them lower scores than the corresponding clean samples. Providing a detailed scoring rubric further enhances this robustness. We further found that fine-tuning an LLM on high-scoring yet biased responses can significantly degrade its performance, highlighting the risk of training on biased data. We also discovered that the judged scores correlate with task difficulty: a challenging dataset like GPQA yields lower average scores, whereas an open-ended reasoning dataset (e.g., JudgeLM-val) sees higher average scores. Finally, we proposed four potential mitigation strategies to ensure fair and reliable AI judging in practical communication scenarios.
@@ I Introduction
Large Language Models (LLMs) are increasingly employed as automated judges to evaluate the quality of AI-generated responses in place of human annotators [1]. This LLM-as-a-Judge paradigm offers scalable and flexible evaluation, providing natural-language feedback and rapid scoring across diverse tasks [2]. The communications and networking industry has begun exploring such AI evaluators in domain-specific applications. For example, LLM-based agents are being considered to assist network operations or customer support, with their responses then judged automatically by another LLM. In fact, generative AI models like ChatGPT are being investigated for intent-based network management and monitoring [3], where an LLM judge could help assess the correctness of an AI assistant’s recommendations before applying them to critical network systems. This promises efficiency gains in telecom operations, but it also raises a pressing question: Can we trust LLM-based evaluators to be fair and accurate in such high-stakes communication scenarios?
@@
Figure 1: Overview of the LLM-as-a-Judge framework in communication systems, illustrating how AI assistants in network management and telecom support are automatically evaluated by an LLM-based judge. We test 11 different types of bias in the experiments.
In this paper, we systematically examine the impact of bias in LLM-as-Judge systems under point-wise settings within communication scenarios, as illustrated in Fig. 1. We focus on how an LLM judge’s scoring is influenced by answers that exhibit specific bias-inducing features. We consider 11 types of biases [6], including subtle implicit biases like verbosity, rich content, or chain-of-thought, as well as more explicit biases like gendered context, authoritative references, or injected factual errors. Our experiments reveal several key findings. First, we found that a well-configured LLM judge, especially one prompted with a detailed evaluation rubric and strict criteria, is largely immune to implicit biases, e.g., verbosity, chain-of-thought, and content richness, and appropriately penalizes explicit biases such as gender and factual error biases. In our tests, none of the bias-injected answers received higher average scores than their clean counterparts under a rigorous judge prompt. In fact, injecting biases significantly reduced the judged score. For example, adding an authoritative reference lowered GPT-Judge’s score from 9.12 to 3.94, and verbosity bias reduced it from 9.12 to 8.78. Second, considering smaller companies often fine-tune local LLMs using high-scoring responses, we assess the downstream impact of training on implicitly biased yet high-scoring data. The results show that fine-tuning with such data significantly degrades model performance, yielding even lower scores than the original pretrained model. Third, we demonstrate that these bias effects generalize across multiple datasets, from knowledge-intensive Q&A (MMLU-Pro [7] and GPQA [8]) to open-ended reasoning tasks (JudgeLM [9]). At the end, we discuss several mitigation strategies for building more trustworthy LLM-as-a-Judge systems, including robust prompt design to enforce objective evaluation criteria, automated bias detection before scoring, model calibration to reduce sensitivity to superficial features, and ensemble judging with human oversight for high-stakes decisions. Our goal is to promote more consistent, unbiased, and trustworthy AI evaluation in communication systems.
@@ II-A LLM-as-a-Judge in Communication Systems
Large language models (LLMs) have recently been repurposed as automated “judges” that evaluate the quality of AI-generated content, often replacing human annotators [4, 10]. In this LLM-as-a-Judge paradigm, a powerful LLM (e.g., GPT-Judge) is prompted to assess responses either by scoring single answers against criteria or by comparing pairs of answers to decide which is better. The appeal of LLM-based evaluation is its scalability and flexibility: a judge model can provide rapid, detailed feedback in natural language and adapt its criteria to different tasks [2]. Compared to fixed metrics or limited human evaluations, LLM judges offer more nuanced and consistent assessments across large datasets. They have demonstrated utility in domains from open-ended Q&A and code generation to conversational agents and beyond.
In the communications and networking field, LLM-as-a-Judge is emerging as a tool for automating evaluations in specialized scenarios. For example, a network operations center might employ an AI assistant to suggest configuration changes or pinpoint faults, and use an LLM judge to automatically score these suggestions before deployment. This provides continuous, real-time vetting of AI-driven decisions in critical infrastructure, essentially adding a safety check for network changes. Likewise, telecom customer-support chatbots can be overseen by an LLM judge that rates the bot’s answers for quality and correctness. Researchers have even explored applying ChatGPT-like models to intent-based network management [11, 12], where an LLM judge could ensure the correctness of an AI’s recommended network actions before they are executed. In all such cases, the LLM judge acts as an arbitration layer, providing an objective-sounding evaluation of AI outputs without direct human intervention. This approach can greatly improve the efficiency and scalability of network management and customer service. However, its reliability is paramount: if the LLM judge is biased or inaccurate, it may lead to misguided decisions in high-stakes communication systems. Thus, understanding and mitigating any unfairness or errors in LLM-as-a-Judge is especially critical for safe adoption in networking applications.
@@ II-B Evaluation Bias in LLM-as-a-Judge
  • Verbosity (Length) Bias. Relatedly, LLM judges often show a direct preference for longer answers. Lengthy responses tend to receive higher scores simply due to their volume or thoroughness of wording, regardless of whether the added text improves correctness or clarity. This “longer is better” bias risks penalizing succinct answers and rewarding verbosity for its own sake.
@@
  • Authority (Reference) Bias. Judges can be swayed by answers that invoke authoritative sources or formal citations, granting them undue credibility. If an answer includes references to research, standards, or expert names (even if irrelevant or fabricated), an LLM judge might score it higher. This bias reflects an over-reliance on authority cues, i.e., the model conflates the presence of references with answer quality.
  • Factual Error (Misinformation) Bias. The judge’s scoring can be insensitive to the correctness of factual content. For example, an answer containing a subtle factual mistake but phrased confidently may still receive a high score, effectively allowing misinformation to pass as acceptable [6]. This oversight means incorrect or misleading answers might escape proper penalization, which is especially risky in network operations or other high-stakes communication scenarios where accuracy is paramount.
@@
  • Gender/Demographic Bias. The evaluation may shift based on the mentioned gender or identity of entities in the answer. For example, if a response casually reveals the responder’s gender or belongs to a certain group, a biased judge might rate it differently (e.g., due to stereotypes or differential expectations). We include tests for gender-based bias as a representative case of diversity bias, noting that answers associated with certain demographics should ideally be judged on content alone.
@@ III Evaluation Bias Impact on LLM-as-a-judge
(a) GPT-Judge
(b) JudgeLM
Figure 2: Average judge scores for answers with $1$ different injected biases.
@@ III-A Experimental Setup
Judge Models and Prompting Strategies. For evaluating the answers, we employed GPT-Judge(prompted on GPT-4o) and JudgeLM(fine-tuned on Vicuna-7B) as judges. We instruct the judge model to rate each answer on a numerical 1–10 scale, with 10 representing the highest level of response quality. We adopt two different prompting strategies: (1) a Detailed Rubric Prompt adapted from the Google Vertex Prompt in JudgeBench [13]. This prompt provides the model with clear criteria and requests a step-by-step reasoning before giving a score, encouraging a thorough and objective evaluation. (2) a Minimal Prompt that simply asks the model, “Please score this answer from 1 to 10” with little additional guidance. This represents a naive usage scenario and tests if the model’s intrinsic biases surface more when unguided.
Evaluation dataset. We first collect a public Q&A dataset from [6], which contains 142 questions spanning various domains and each came with 4 GPT-generated bias answers and 1 clean answer. We then used GPT-Judge as a judge to score every answer on a 1–10 scale. From this set, we select the top 50 unique questions with the highest GPT-Judge scores (typically 9–10) and their corresponding clean answers.
@@ III-B Performance Evaluation
Impact of Bias Type. We first assessed each bias type individually, using the 50 biased responses generated for that specific bias across all 50 questions. Fig. 2 summarizes the average scores (on a 1–10 scale) assigned by the GPT-Judge judge to answers containing each bias. As shown in the Fig. 2, there are clear disparities in how different biases affect the judged quality of answers.
In general, all biased answers received lower mean scores than their corresponding clean (unbiased) answers, though the degree of score reduction varied widely by bias type. Introducing a factual error bias had the most severe effect, with incorrect answers averaging only 4.98 for JudgeLM. In contrast, adding implicit bias like content-richness or verbosity had minimal impact, with average scores remaining relatively high at 8.82 and 8.80, respectively. These results highlight that certain biases can strongly skew an LLM judge’s scoring. A likely reason is that some injected biases directly undermine the clarity or correctness of an answer from the judge’s perspective.
Interestingly, in our experiments, the verbose and chain-of-thought reasoning answers actually scored lower than the concise clean answers. One explanation is that our evaluation prompt and criteria encouraged focus on correctness and relevance, so unnecessarily long or self-reasoning answers were viewed as less clear. This indicates that GPT-Judge, when properly prompted, does not always reward length for its own sake, in fact, overly verbose answers were treated as inferior in quality.
@@
Fine-tuning Dataset Sample Num GPT-Judge Score
@@
Specifically, we construct the training dataset from the Evaluation dataset. The “Clean” dataset contains unbiased, high-quality answers. The “Bias” dataset includes top-scoring bias-injected answers (one per question) for the same 50 questions. In addition, we evaluate fine-tuning using a larger, generic instruction-following dataset, i.e., Alpaca [14], from which we sample 1,000 instances. Variants such as “Alpaca + Clean” and “Alpaca + Bias” indicate that 50 specialized samples were added to the 1k Alpaca subset. “Alpaca + Raw” denotes that all samples are drawn exclusively from the Alpaca dataset.
@@
Fine-tuning Dataset Sample Num JudgeLM Score
@@
TABLE III: Average GPT-Judge evaluation scores for responses produced by Llama 3.1-8B-Instruct models trained on different datasets, under two evaluation prompt styles: the detailed Elaborated prompt and the minimal Easy prompt.
Evaluation Model MMLU-Pro GPQA JudgeLM val
Elaborated Easy Elaborated Easy Elaborated Easy
Bias 50 6.60 7.04 4.62 5.22 6.80 7.78
Clean 50 6.84 7.45 5.43 5.64 7.97 8.18
Alpaca+Bias (1k+50) 6.86 7.23 4.45 5.20 7.71 8.16
Alpaca+Clean (1k+50) 6.85 7.30 4.92 5.16 7.56 8.18
Alpaca+Raw (1k+50) 6.73 7.07 4.61 5.18 7.90 8.14
@@
Impact of different test datasets. Apart from the basic datasets used in JudgeBench [6], we further evaluated on three diverse benchmarks, i.e., MMLU-Pro [7], GPQA [8], and the JudgeLM [9], with more difficulty.
@@
MMLU-Pro extends the original Massive Multitask Language Understanding benchmark with more reasoning-intensive questions and expanded multiple-choice options (from 4 to 10), removing trivial items to better test LLMs’ reasoning and knowledge.
@@
GPQA is a challenging “Google-proof” Q&A dataset of 448 graduate-level science questions (e.g., biology, physics, chemistry), designed to prevent simple lookup. It consists of 448 expert-written multiple-choice questions, and even domain experts with PhDs only achieve about 65% accuracy on this benchmark, underscoring its toughness.
@@
JudgeLM validation set [9] includes open-ended evaluation tasks. Unlike MMLU-Pro or GPQA, it does not have single correct answers but instead assesses response quality in subjective, free-form contexts such as conversational or explanatory outputs.
@@
As shown in Table LABEL:tab:prompt-impact, model scores from the GPT-Judge judge decrease as task difficulty increases. GPQA produces the lowest average scores, reflecting the challenge posed by its graduate-level questions, whereas the JudgeLM validation tasks achieve the highest scores (around 7–8 on average) due to they are more open-ended and subjective. Models fine-tuned on high-quality clean answers achieved the best scores on all benchmarks, whereas those fine-tuned on bias-injected answers (Bias-50) performed worse overall. For instance, on the GPQA set with the Easy prompt, the Clean-50 model attained an average GPT-Judge score around 5.6, noticeably higher than the Bias-50 model’s 5.2.
@@
The trend is clear: using a detailed, structured prompt leads to systematically lower scores (harsher judgments), while the minimal prompt yields higher scores (more lenient evaluations) for the same answers. For example, the LLaMA 3.1 evaluation model trained on 50 bias samples (Bias-50) obtained an average score of 7.04 on MMLU-Pro questions under the detailed rubric prompt, compared to 6.6 under the minimal prompt. A similar trend is observed across other datasets and evaluation models.
A plausible explanation is that the Elaborated prompt forces the GPT-Judge judge to follow a step-by-step evaluation rubric, which in turn reduces the influence of superficial answer features. This structured reasoning likely filters out spurious advantages such as excessive length, authoritative-sounding references, or emotional tone that might otherwise sway the judge. By focusing the evaluation on factual correctness, relevance, and adherence to instructions, the detailed prompt yields more objective (albeit stricter) scoring. In contrast, the Minimal prompt provides little guidance, so the judge model may unconsciously reward those superficial features, leading to higher, but less reliable, scores for the same content.

### v3

@@ Abstract
Large Language Models (LLMs) are increasingly being used to autonomously evaluate the quality of content in communication systems, e.g., to assess responses in telecom customer support chatbots. However, the impartiality of these AI “judges” is not guaranteed, and any biases in their evaluation criteria could skew outcomes and undermine user trust. In this paper, we systematically investigate judgment biases across 6 LLM-as-a-judge models spanning both prompt-based and finetuned judges under the pointwise scoring setting, encompassing 11 types of biases that cover both implicit and explicit forms. We observed that state-of-the-art LLM judges demonstrate robustness to biased inputs, generally assigning them lower scores than the corresponding clean samples. We further found that fine-tuning an LLM on high-scoring yet biased responses can significantly degrade its performance, highlighting the risk of training on biased data. We also discovered that the judged scores correlate with task difficulty: a challenging dataset like GPQA yields lower average scores, whereas an open-ended reasoning dataset (e.g., JudgeLM-val) sees higher average scores. Finally, we proposed four potential mitigation strategies to ensure fair and reliable AI judging in practical communication scenarios.
@@ I Introduction
Large Language Models (LLMs) are increasingly used as automated judges to evaluate AI-generated responses, offering scalable assessment with natural-language feedback and rapid scoring [1, 2]. In parallel, transformer-based architectures have also been applied to sequential prediction in physical systems, forecasting dynamical time series from input histories or initial conditions [3, 4]. Nowadays, the communications and networking industry is beginning to adopt LLM-based evaluators in domain-specific workflows. For example, LLM agents are being explored for network operations and customer support, where another LLM can automatically judge the correctness of an assistant’s recommendations before deployment. While this promises efficiency gains, it also raises a pressing question: Can we trust LLM-based evaluators to be fair and accurate in high-stakes communication scenarios?
@@
Figure 1: Overview of the LLM-as-a-Judge framework in communication systems, illustrating how AI assistants in network management and telecom support are automatically evaluated by an LLM-based judge.
In this paper, we systematically examine the impact of bias in LLM-as-Judge systems under pointwise settings within communication scenarios, as illustrated in Fig. 1. We focus on how an LLM judge’s scoring is influenced by answers that exhibit specific bias-inducing features. We consider 11 types of biases [7], including subtle implicit biases like verbosity, rich content, or chain-of-thought, as well as more explicit biases like gendered context, authoritative references, or injected factual errors. Our experiments reveal several key findings. First, we find that a well-configured LLM judge can penalize biased inputs by scoring them lower than clean counterparts, and it is more sensitive to explicit biases (e.g., authority) than implicit ones (e.g., verbosity). For instance, an incorrect/irrelevant authoritative reference drops GPT-Judge’s score from 9.12 to 3.94, while verbosity reduces it only from 9.12 to 8.78. Furthermore, we observe that although detailed prompts lower absolute scores, the relative penalty on biased answers stays stable across prompting strategies, suggesting bias sensitivity is intrinsic to LLM judges rather than prompt-induced. Second, considering smaller companies often fine-tune local LLMs using high-scoring responses, we assess the downstream impact of training on implicitly biased yet high-scoring data. Our results show that although such fine-tuning can improve performance over the pretrained model, models fine-tuned on biased data consistently underperform those fine-tuned on clean data under the LLM-as-a-Judge evaluation framework. Third, we demonstrate that these bias effects generalize across multiple datasets, from knowledge-intensive Q&A (MMLU-Pro [8] and GPQA [9]) to open-ended reasoning tasks (JudgeLM [10]). At the end, we discuss several mitigation strategies for building more trustworthy LLM-as-a-Judge systems, including robust prompt design to enforce objective evaluation criteria, automated bias detection before scoring, model calibration to reduce sensitivity to superficial features, and ensemble judging with human oversight for high-stakes decisions. Our goal is to enable trustworthy LLM-as-a-judge evaluation of AI generated content in communication systems, by making the judging process more consistent and fair (less sensitive to superficial bias-inducing features and more responsive to explicit errors) through a reproducible pointwise evaluation protocol and actionable mitigation guidelines.
@@ II-A LLM-as-a-Judge in Communication Systems
Large language models (LLMs) have recently been repurposed as automated “judges” that evaluate the quality of AI-generated content, often replacing human annotators [5]. In this LLM-as-a-Judge paradigm, a powerful LLM (e.g., GPT-4o) is prompted to assess responses either by scoring single answers against criteria or by comparing pairs of answers to decide which is better. The appeal of LLM-based evaluation is its scalability and flexibility: a judge model can provide rapid, detailed feedback in natural language and adapt its criteria to different tasks [2]. Compared to fixed metrics or limited human evaluations, LLM judges offer more nuanced assessments and demonstrate stable evaluation trends when applied at scale across large datasets. They have demonstrated utility in domains from open-ended Q&A and code generation to conversational agents and beyond.
In the communications and networking field, LLM-as-a-Judge is emerging as a tool for automating evaluations in specialized scenarios. For example, a network operations center might employ an AI assistant to suggest configuration changes or pinpoint faults, and use an LLM judge to automatically score these suggestions before deployment. This provides continuous, real-time vetting of AI-driven decisions in critical infrastructure, essentially adding a safety check for network changes. Likewise, telecom customer-support chatbots can be overseen by an LLM judge that rates the bot’s answers for quality and correctness. Researchers have even explored applying ChatGPT-like models to intent-based network management [11], where an LLM judge could ensure the correctness of an AI’s recommended network actions before they are executed. In all such cases, the LLM judge acts as an arbitration layer, providing an objective-sounding evaluation of AI outputs without direct human intervention. This approach can greatly improve the efficiency and scalability of network management and customer service. However, its reliability is paramount: if the LLM judge is biased or inaccurate, it may lead to misguided decisions in high-stakes communication systems. Thus, understanding and mitigating any unfairness or errors in LLM-as-a-Judge is especially critical for safe adoption in networking applications.
Beyond LLM-as-a-Judge, recent work explores integrating generative models into networking systems for management tasks (e.g., traffic forecasting and policy generation) [12]. These studies primarily evaluate task-level performance and system pipelines, whereas our paper examines the evaluation layer itself.
@@ II-B Evaluation Bias in LLM-as-a-Judge
  • Verbosity (Length) Bias. Relatedly, LLM judges often show a direct preference for longer answers. Lengthy responses tend to receive higher scores simply due to their volume or thoroughness of wording, regardless of whether the added text improves correctness or clarity. This “longer is better” bias risks penalizing succinct answers and rewarding verbosity for its own sake. In a Network Operations Center (NOC), an LLM judge with such bias might favor a lengthy, vague troubleshooting guide over a concise, precise CLI command that resolves the fault immediately, leading to operational inefficiencies.
@@
  • Authority (Reference) Bias. Judges can be swayed by answers that invoke authoritative sources or formal citations, granting them undue credibility. If an answer includes references to research, standards, or expert names (even if irrelevant or fabricated), an LLM judge might score it higher. In telecom standard implementation, a response that hallucinates non-existent 3GPP or IEEE citations might be scored higher than a correct answer without citations, potentially misleading engineers during protocol configuration.
  • Factual Error (Misinformation) Bias. The judge’s scoring can be insensitive to the correctness of factual content. For example, an answer containing a subtle factual mistake but phrased confidently may still receive a high score, effectively allowing misinformation to pass as acceptable [7]. This oversight means incorrect or misleading answers might escape proper penalization, which is especially risky in network operations or other high-stakes communication scenarios where a confident but factually incorrect configuration parameter (e.g., wrong IP subnet mask or BGP weight) can lead to widespread service outages.
@@
  • Gender Bias. The evaluation may shift based on the mentioned gender or identity of entities in the answer. For instance, in telecom customer support analysis, a biased judge might undervalue technical feedback from female customers while treating identical reports from male customers as critical system faults. We include tests for gender-based bias as a representative case of diversity bias, noting that answers associated with certain demographics should ideally be judged on content alone.
@@ III Evaluation Bias Impact on LLM-as-a-judge
Figure 2: Average judge scores for answers with $1$ different injected biases evaluated by (a) prompt-based judges and (b) fine-tuned judges.
@@ III-A Experimental Setup
Judge Models and Prompting Strategies. We conducted experiments on 6 representative LLM-as-a-judge models, spanning both prompt-based judges and fine-tuned open-source judges.
  • Prompt-based Judges. We evaluate GPT-4o, GPT-5.1, and Claude Sonnet 4.5, 3 state-of-the-art proprietary foundation models from OpenAI and Anthropic with strong reasoning and evaluation capabilities across diverse benchmarks.
  • Fine-tuned Judges. We consider JudgeLM [10], Con-J-7B-Instruct [13], and CompassJudger-2-7B-Instruct [14], 3 open-source judges explicitly fine-tuned for evaluation tasks, built on Vicuna-7B, Qwen-7B, and Qwen2.5-7B-Instruct, respectively.
These models span diverse architectures, and we instruct each judge to score responses on a 1–10 scale, where 10 denotes the highest quality. We adopt two different prompting strategies: (1) a Detailed Rubric Prompt adapted from the Google Vertex Prompt in JudgeBench [15]. This prompt provides the model with clear criteria and requests a step-by-step reasoning before giving a score, encouraging a thorough and objective evaluation. (2) a Minimal Prompt that simply asks the model, “Please score this answer from 1 to 10” with little additional guidance. This represents a naive usage scenario and tests if the model’s intrinsic biases surface more when unguided.
Evaluation dataset. We first collect a public Q&A dataset from [7], which contains 142 questions spanning various domains and each came with 4 GPT-generated bias answers and 1 clean answer. We then used GPT-Judge (prompted on GPT-4o) as a judge to score every answer on a 1–10 scale. From this set, we select the top 50 unique questions with the highest GPT-Judge scores (typically 9–10) and their corresponding clean answers.
@@ III-B Performance Evaluation
Impact of Bias Type. We first evaluate each bias independently using 50 biased responses and their clean counterparts across 50 questions, reducing instance-level variance and improving statistical robustness. Fig. 2 reports the average scores assigned by all 6 judge models to responses with each bias under the Detailed Rubric Prompt.
There are clear disparities in how different biases affect the judged quality of answers. In general, almost all biased answers received lower mean scores than their corresponding clean (unbiased) answers, though the degree of score reduction varied widely by bias type. Factual error bias cause one of the sharpest score drops across judge architectures. For example, reducing scores from 9.12 to 4.00 for GPT-4o, from 8.96 to 3.22 for Claude Sonnet 4.5, and from 9.20 to 3.38 for GPT-5.1. In contrast, adding implicit bias like content-richness or verbosity had minimal impact, with average scores remaining relatively high at 8.82 and 8.80, respectively for JudgeLM. These results highlight that certain biases can strongly skew an LLM judge’s scoring. A likely reason is that these injected biases directly undermine the clarity or relevance of an answer from the judge’s perspective.
Interestingly, verbose and chain-of-thought responses are usually scored slightly lower than clean answers across most judges, though the magnitude varies by model. A likely reason is that our rubric emphasizes correctness and relevance, making unnecessarily long or self-reasoning responses appear less clear. This suggests that, when properly prompted, modern LLM judges can distinguish informative content from superficial verbosity rather than consistently favoring longer answers.
@@
Fine-tuning Dataset No. of Samples GPT-Judge Score
@@
Specifically, we construct the training dataset from the Evaluation dataset. The “Clean” dataset contains unbiased, high-quality answers. The “Bias” dataset includes top-scoring bias-injected answers (one per question) for the same 50 questions. In addition, we evaluate fine-tuning using a larger, generic instruction-following dataset, i.e., Alpaca, from which we sample 1,000 instances. Variants such as “Alpaca + Clean” and “Alpaca + Bias” indicate that 50 specialized samples were added to the 1k Alpaca subset. “Alpaca + Raw” denotes that all samples are drawn exclusively from the Alpaca dataset.
@@
Fine-tuning Dataset No. of Samples JudgeLM Score
@@
TABLE III: Average evaluation scores for responses produced by Llama 3.1-8B-Instruct models trained on different datasets, evaluated by GPT-Judge and JudgeLM under two prompt styles. Changes relative to Easy prompt are shown in parentheses.
Judge Model MMLU-Pro GPQA JudgeLM val
Minimal Elaborate Minimal Elaborate Minimal Elaborate
GPT-Judge Bias 50 7.04 6.60 ($\downarrow$0.44) 5.22 4.62 ($\downarrow$0.60) 7.78 6.80 ($\downarrow$0.98)
Clean 50 7.45 6.84 ($\downarrow$0.61) 5.64 5.43 ($\downarrow$0.21) 8.18 7.97 ($\downarrow$0.21)
Alpaca+Bias (1k+50) 7.23 6.86 ($\downarrow$0.37) 5.20 4.45 ($\downarrow$0.75) 8.16 7.71 ($\downarrow$0.45)
Alpaca+Clean (1k+50) 7.30 6.85 ($\downarrow$0.45) 5.16 4.92 ($\downarrow$0.24) 8.18 7.56 ($\downarrow$0.62)
Alpaca+Raw (1k+50) 7.07 6.73 ($\downarrow$0.34) 5.18 4.61 ($\downarrow$0.57) 8.14 7.90 ($\downarrow$0.24)
JudgeLM Bias 50 4.95 5.07 ($\uparrow$0.12) 3.56 3.42 ($\downarrow$0.14) 8.18 7.96 ($\downarrow$0.22)
Clean 50 5.23 5.34 ($\uparrow$0.11) 4.09 3.91 ($\downarrow$0.18) 8.11 8.10 ($\downarrow$0.01)
Alpaca+Bias (1k+50) 5.14 5.17 ($\uparrow$0.03) 3.42 3.63 ($\uparrow$0.21) 7.43 7.32 ($\downarrow$0.11)
Alpaca+Clean (1k+50) 5.15 5.19 ($\uparrow$0.04) 3.70 3.63 ($\downarrow$0.07) 7.42 7.37 ($\downarrow$0.05)
Alpaca+Raw (1k+50) 5.20 5.10 ($\downarrow$0.10) 3.32 3.43 ($\uparrow$0.11) 7.07 6.85 ($\downarrow$0.22)
@@
Impact of different test datasets. Apart from public Q&A dataset constructed in [7], we further evaluated on three diverse benchmarks, i.e., MMLU-Pro [8], GPQA [9], and JudgeLM [10], which were strategically selected to simulate the multifaceted scenarios encountered in communication systems.
@@
MMLU-Pro extends the original Massive Multitask Language Understanding benchmark with more reasoning-intensive questions and expanded multiple-choice options (from 4 to 10), removing trivial items to better test LLMs’ reasoning and knowledge. This serves as a proxy for expert-level network troubleshooting that involves high-difficulty reasoning steps, where the judge must logically validate solutions among multiple complex possibilities.
@@
GPQA is a challenging “Google-proof” Q&A dataset of 448 graduate-level science questions (e.g., biology, physics, chemistry), designed to prevent simple lookup. It consists of 448 expert-written multiple-choice questions, and even domain experts with PhDs only achieve about 65% accuracy on this benchmark, underscoring its toughness. We utilize this dataset to simulate high-stakes technical support, where the judge must discern subtle errors in complex protocols akin to core network diagnostics.
@@
JudgeLM validation set [10] includes open-ended evaluation tasks. Unlike MMLU-Pro or GPQA, it does not have single correct answers but instead assesses response quality in subjective, free-form contexts such as conversational or explanatory outputs. This aligns with scenarios like customer service systems, which frequently involve open-ended tasks requiring the judge to evaluate the relevance and helpfulness of non-standardized responses.
@@
As shown in Table LABEL:tab:prompt-impact, model scores from both GPT-Judge and JudgeLM decrease as task difficulty increases. GPQA produces the lowest average scores, reflecting the challenge posed by its graduate-level questions, whereas the JudgeLM validation tasks achieve the highest scores (around 7–8 on average) due to they are more open-ended and subjective. Across both GPT-Judge and JudgeLM evaluators, models fine-tuned on high-quality clean answers achieved the best scores on all benchmarks, whereas those fine-tuned on bias-injected answers (Bias-50) performed worse overall. For instance, on the GPQA set with the Minimal prompt, the Clean-50 model attained an average GPT-Judge score of 5.64, noticeably higher than the Bias-50 model’s 5.22.
@@
The trend is clear for GPT-Judge. Using a detailed, structured prompt leads to systematically lower scores (harsher judgments), while the minimal prompt yields higher scores (more lenient evaluations) for the same answers. For example, the LLaMA 3.1 evaluation model trained on 50 bias samples (Bias-50) obtained an average score of 7.04 on MMLU-Pro questions under the minimal prompt, compared to 6.6 under the detailed rubric prompt. A similar trend is observed across other datasets and evaluation models. In contrast, JudgeLM exhibits lower sensitivity to prompt variations, with its scores remaining relatively stable across different prompt styles. While minor fluctuations are observed, the overall scoring trends remain consistent under both the Minimal and Elaborate prompts.
A plausible explanation is that the Elaborated prompt forces the GPT-Judge judge to follow a step-by-step evaluation rubric, which in turn reduces the influence of superficial answer features. This structured reasoning likely filters out spurious advantages such as excessive length, authoritative-sounding references, or emotional tone that might otherwise sway the judge. By focusing the evaluation on factual correctness, relevance, and adherence to instructions, the detailed prompt yields more objective (albeit stricter) scoring. In contrast, JudgeLM’s stability likely stems from being fine-tuned as a dedicated evaluator on large-scale annotated scoring data, allowing it to internalize the evaluation criteria during training. Consequently, JudgeLM’s scoring decisions rely more on its learned evaluation function and are less dependent on external prompt guidance.